# Efficient Transformed Gaussian Process State-Space Models for Non-Stationary High-Dimensional Dynamical Systems

### v2

@@ III Efficient Transformed Gaussian Process State-Space Models
We first introduce the proposed efficient and flexible modeling framework for high-dimensional state spaces in Section III-A. Section III-B then presents the corresponding Bayesian variational inference algorithm. Finally, the details of EnKF-aided variational lower bound evaluation are provided in Section III-C.
@@ III-A Efficient Modeling for High-Dimensional State Space
Figure 2: (2(a)) Two warped GPs, see Eq. (10). $f_{1}(\cdot)$ effectively fits the data, while $f_{2}(\cdot)$ falls short. (2(b)) Independent GPs ($\tilde{f}_{1}(\cdot)$, $\tilde{f}_{2}(\cdot)$) vs our proposed ETGP (${f}_{1}(\cdot)$, ${f}_{2}(\cdot)$) from Fig. 3. Both methods capture the underlying function reasonably well, though the GP $\tilde{f}_{1}(\cdot)$ exhibits slightly worse fit and overestimated uncertainty.
In Section II-C, we discussed the computational and modeling challenges associated with using separate GPs for each dimension of the transition function in high-dimensional state spaces. To address these issues, in [38], we proposed to warp a single shared GP using normalizing flows [43] to fit each dimension of the transition function, that is,
@@
where each dimension-specific normalizing flow, ${{\mathbb{G}}}_{\bm{\theta}_{d}}(\cdot)$, is an element-wise, invertible and differentiable (bijective) function parameterized by $\bm{\theta}_{d}\in\Theta$. Two specific examples are provided below.
@@
The primary advantage of this methodology lies in its elimination of multiple independent GPs by leveraging computationally efficient flow transformations, thereby enabling circumvent the $\mathcal{O}(d_{x}M^{3})$ computational complexity and quadratic parameter proliferation mentioned in Section II-C. Additionally, the warped GPs are dependent which could potentially alleviate the limitation of independence modeling in conventional GPSSMs.
However, the warped GP in Eq. (10) remains stationary and, being based on a single GP, may still lack the flexibility required to capture complex latent state dynamics, leading to suboptimal learning performance [38]. An intuitive illustrative example is shown in Fig. 2(a), where the learned element-wise flow $\mathbb{G}_{\bm{\theta}_{1}}$ successfully warps $\tilde{f}$ to obtain $f_{1}$, achieving accurate fitting of the underlying function’s simple pattern. However, when faced with more complex, non-stationary data trajectories, the warped GP, $f_{2}$, struggles to capture these patterns, resulting in huge modeling uncertainty. For a more comprehensive comparison, Fig. 2(b) also presents the fitting results obtained using two independent GPs and the ETGP model introduced later in this paper. Both methods effectively capture the underlying functions, although the GP $\tilde{f}_{1}(\cdot)$ demonstrates a slightly inferior fit and inflated uncertainty estimates, due to its inherent stationarity.
To address the limitations in modeling flexibility exhibited by existing methods, we draw inspiration from recent advances in adaptive parameterization [44, 47] and introduce non-stationarity and input dependence into the warped GPs. Specifically, we generate non-stationary processes $\{f_{d}(\cdot)\}_{d=1}^{d_{x}}$ by making the flow parameters depend on the input through a shallow neural network, denoted as $\textsc{nn}:\mathbb{R}^{d_{x}}\mapsto\Theta$ and parameterized by a weight vector $\mathbf{w}$. Formally, in the context of SSM, this is expressed as [^2] Note that ${\mathbf{x}}_{t}$ can always be extended to accommodate control systems incorporating a deterministic control input $\bm{c}_{t}\in\mathbb{R}^{d_{c}}$ by augmenting the latent state to $[{\mathbf{x}}_{t}^{\top},\bm{c}_{t}^{\top}]^{\top}\in\mathbb{R}^{d_{x}+d_{c}}$. For brevity, however, we omit explicit reference to $\bm{c}_{t}$ in our notation throughout this paper.:
@@
where ${{\bm{\theta}}}_{t}=\{{{\bm{\theta}}}_{t,d}\}_{d=1}^{d_{x}}$. In this case, the flow parameters vary with the input at each time step $t$, enabling the capture of local variation patterns and facilitating the modeling of time-varying transition functions. Additionally, note that the parameter amortization learning shown in Eq. (11) can avoid that the number of flow parameters increases linearly with time step $t$. Moreover, sharing neural network parameters across flow parameterizations further improves modeling efficiency.
Figure 3: ETGP transition function in SSMs
@@
It is noteworthy that this modeling approach, also known as the efficient transformed Gaussian process (ETGP) in the GP literature, was initially proposed to address large-scale multi-class classification problems [48]. The efficiency stems from the use of a single GP rather than multiple GPs for multi-class settings, thereby ensuring that the computational cost of GP operations remains constant with respect to the number of classes. This leads to a substantial reduction in overall computations. In this work, we adapt it to model the high-dimensional transition function in SSMs. Additionally, ETGP implicitly characterizes $d_{x}$ dependent random processes [42], a dependence that has been overlooked in prior GPSSM literature. A more detailed insight is provided below.
@@
This result can be built upon the Gaussianity and the linearity property of covariance, which can be found in Appendix A. ∎
With the computationally efficient and flexible function prior defined by ETGP $\bm{f}(\cdot)\!=\!\{f_{d}(\cdot)\}_{d=1}^{d_{x}}$, we can define our proposed SSM, termed ETGPSSM, compactly expressed as follows:
@@
where the ETGP prior is characterized by the associated GP $\tilde{f}(\cdot)$, neural network weight prior $p_{\bm{\psi}}(\mathbf{w})$, and the normalizing flows, see Eq. (12). In this paper, unless otherwise specified, we focus on the linear flow shown in Example 3 for simplicity and illustrative purposes. Additionally, the linear flow requires learning only $2$ parameters per dimension, making it more efficient compared to other types of normalizing flows. We also assume that the prior distribution of the initial state, $p({\mathbf{x}}_{0})$, is Gaussian and known for simplicity.
@@
The primary objective in ETGPSSM, as in GPSSM, is the joint estimation of model hyperparameters $\bm{\vartheta}\!=\!\{{{\bm{\theta}}}_{gp},\bm{\psi},\mathbf{Q},\mathbf{R}\}$ and inference of both the ETGP and latent states. More specifically, given a sequence of observations ${\mathbf{y}}_{1:T}=\{{\mathbf{y}}_{t}\}_{t=1}^{T}$, with $T\in\mathbb{Z}_{+}$, the task is to estimate $\bm{\vartheta}$ and compute the posterior distributions $p(\bm{f}(\cdot)|{\mathbf{y}}_{1:T})$ for the ETGP and $p({\mathbf{x}}_{0:T}|{\mathbf{y}}_{1:T})$ for the latent states, where ${\mathbf{x}}_{0:T}=\{{\mathbf{x}}_{t}\}_{t=0}^{T}$.
However, the marginal likelihood $p(\mathbf{y}_{1:T}|\bm{\vartheta})$ is generally intractable, rendering both hyperparameter estimation via model evidence maximization and posterior inference computationally infeasible. Furthermore, unlike standard GPs that admit closed-form expressions for the prior and posterior distributions (see Eqs. (2) and (3)), the ETGP constitutes an implicit stochastic process. Consequently, the marginal prior [^3] Obtaining $p(\bm{f}(\cdot))$ by marginalizing over the GP and $\mathbf{w}$ is generally intractable. and posterior distributions are analytically intractable [42], posing significant challenges for inference. The next subsections introduce a variational inference framework that enables joint learning of model hyperparameters and Bayesian inference of the ETGP and latent states.
@@ III-B Approximate Bayesian Inference
To address the intractability of the marginal distribution $p({\mathbf{y}}_{1:T}|\bm{\vartheta})$, we resort to variational inference, which involves maximizing an evidence lower bound (ELBO) on the logarithm of the marginal likelihood [49]. The ELBO, denoted as $\mathcal{L}$, can be generally constructed as follows:
$\mathcal{L}=\mathbb{E}_{q({\mathbf{x}}_{0:T},\bm{f}(\cdot))}\left[\log\frac{p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0:T},\bm{f}(\cdot))}{q({\mathbf{x}}_{0:T},\bm{f}(\cdot))}\right],$ (17)
so that the difference between $\log p({\mathbf{y}}_{1:T}|\bm{\vartheta})$ and $\mathcal{L}$ equals the Kullback-Leibler (KL) divergence between the introduced variational approximation $q({\mathbf{x}}_{0:T},\bm{f}(\cdot))$ and the true posterior $p({\mathbf{x}}_{0:T},\bm{f}(\cdot)|{\mathbf{y}}_{1:T})$. We then can jointly maximize $\mathcal{L}$ with respect to $\bm{\vartheta}$ and the variational distribution $q({\mathbf{x}}_{0:T},\bm{f}(\cdot))$ to obtain the model hyperparameter estimates and the approximation of the posterior distribution $p({\mathbf{x}}_{0:T},\bm{f}(\cdot)|{\mathbf{y}}_{1:T})$.
@@
However, as noted earlier, the absence of closed-form expressions for the ETGP’s prior and posterior distributions prevents direct expansion of the ELBO in Eq. (17). To address this challenge, rather than operating in the function space of the ETGP, we instead follow its generative construction—illustrated in Fig. 3—and infer the posterior distributions over $\tilde{f}$ and $\mathbf{w}$, which collectively characterize the ETGP posterior. Specifically, by explicitly representing $\tilde{f}$ and $\mathbf{w}$, the joint distribution of the proposed ETGPSSM is given by:
$p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0:T},\mathbf{w},\tilde{f})=p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0:T}|\mathbf{w},\tilde{f})p_{\bm{\psi}}(\mathbf{w})p(\tilde{f}),$ (18)
where $p(\tilde{f})$ represents the GP prior with a bit of notation abuse [31] and the prior of $\mathbf{w}$ is assumed to be Gaussian, $p_{\bm{\psi}}(\mathbf{w})={\cal N}(\mathbf{w}|\bm{0},\operatorname{diag}(\bm{\psi}))$, with diagonal variance. The term $p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0:T}|\mathbf{w},\tilde{f})$ is factorized as:
where the transition term is specified as:

### v3

@@ III Efficient Transformed Gaussian Process State-Space Models
We first present our proposed efficient and flexible modeling framework for high-dimensional state spaces in Section III-A. Next, Section III-B introduces the corresponding EnKF-aided variational inference algorithm. Finally, Section III-C details the evaluation of the EnKF-aided variational lower bound.
@@ III-A Efficient Modeling for High-Dimensional State Space
Figure 2: (2(a)) Two warped GPs, see Eq. (10). $f_{1}(\cdot)$ effectively fits the data, while $f_{2}(\cdot)$ falls short. (2(b)) Independent GPs ($\tilde{f}_{1}(\cdot)$, $\tilde{f}_{2}(\cdot)$) vs our proposed ETGP (${f}_{1}(\cdot)$, ${f}_{2}(\cdot)$) depicted in Fig. 3. Both methods capture the underlying function reasonably well, though the GP $\tilde{f}_{1}(\cdot)$ exhibits slightly worse fit and overestimated uncertainty.
In Section II-C, we discussed the computational and modeling challenges that arise from using separate GPs for each output dimension of the transition function in high-dimensional state spaces. To address these limitations, [38] proposed warping a single shared GP using normalizing flows [44] to model each output dimension of the transition function:
@@
where each dimension-specific normalizing flow, ${{\mathbb{G}}}_{\bm{\theta}_{d}}(\cdot)$, is an element-wise, bijective, and differentiable transformation parameterized by $\bm{\theta}_{d}\in\Theta$. Two concrete examples are presented below.
@@
The main advantage of this approach lies in eliminating the need for multiple independent GPs by using computationally efficient flow transformations. This design bypasses the $\mathcal{O}(d_{x}M^{3})$ computational complexity and quadratic parameter growth discussed in Section II-C. Moreover, since all dimensions share the same underlying GP, the resulting warped functions are inherently dependent, potentially mitigating the limitations of modeling independent outputs in standard GPSSMs.
However, the warped GP in Eq. (10) remains stationary; Also, being derived from a single shared GP, it may lack the expressiveness needed to capture complex, time-varying latent dynamics—resulting in suboptimal performance in practice [38]. An illustrative example is shown in Fig. 2(a), where the learned element-wise flow $\mathbb{G}_{\bm{\theta}_{1}}$ successfully warps $\tilde{f}$ to fit a simple target function $f_{1}$. In contrast, the warped function $f_{2}$ struggles to capture the time-varying and intricate structure, resulting in high uncertainty. For a more comprehensive comparison, Fig. 2(b) shows the results from both two independent GPs and the ETGP model introduced later in this work. While both capture the underlying structure, the stationary GP $\tilde{f}_{1}(\cdot)$ yields inferior fits and inflated uncertainty estimates.
To overcome these limitations in expressiveness, we draw inspiration from recent advances in adaptive parameterization [45, 48] and introduce non-stationarity and input dependence into the warped GPs. Specifically, we generate non-stationary processes $\{f_{d}(\cdot)\}_{d=1}^{d_{x}}$ by conditioning the flow parameters on the input through a shallow neural network $\textsc{nn}:\mathbb{R}^{d_{x}}\mapsto\Theta$ parameterized by weights $\mathbf{w}$. In the SSM context, this is formulated as [^3] Note that ${\mathbf{x}}_{t}$ can always be extended to accommodate control systems incorporating a deterministic control input $\bm{c}_{t}\in\mathbb{R}^{d_{c}}$ by augmenting the latent state to $[{\mathbf{x}}_{t}^{\top},\bm{c}_{t}^{\top}]^{\top}\in\mathbb{R}^{d_{x}+d_{c}}$. For brevity, however, we omit explicit reference to $\bm{c}_{t}$ in our notation throughout this paper.:
@@
where ${{\bm{\theta}}}_{t}=\{{{\bm{\theta}}}_{t,d}\}_{d=1}^{d_{x}}$. This formulation allows the flow parameters to vary with input at each time step $t$, enabling the model to capture local variations and time-varying dynamics. Additionally, note that the parameter amortization learning shown in Eq. (11) prevents the number of flow parameters from growing with time step $t$ while also improving parameter efficiency through weight sharing across dimensions.
@@
Figure 3: ETGP transition function in SSMs
It is noteworthy that this modeling strategy, referred to as the efficient transformed Gaussian process (ETGP) in the GP literature, was originally proposed to tackle large-scale multi-class classification problems [42]. Its efficiency arises from employing a single GP instead of one per class, ensuring that the computational cost of GP operations remains independent of the number of classes. This results in a substantial reduction in overall computation. In this work, we adapt the ETGP framework to model high-dimensional transition functions in SSMs. Notably, ETGP implicitly models $d_{x}$ dependent random processes [43], introducing output correlations that are often neglected in existing GPSSM formulations. We elaborate further on this below.
@@
This result follows from the Gaussianity of the underlying process and the linearity of covariance, as detailed in Appendix C. ∎
With the computationally efficient and flexible function prior defined by ETGP $\bm{f}(\cdot)\!=\!\{f_{d}(\cdot)\}_{d=1}^{d_{x}}$, we define our proposed SSM, termed ETGPSSM, as follows:
@@
where the ETGP prior is defined by the base GP $\tilde{f}(\cdot)$, neural network weight prior $p_{\bm{\psi}}(\mathbf{w})$, and the normalizing flows, see Eq. (12). Unless stated otherwise, in this work we focus on the linear flow described in Example 3 for simplicity and illustration. This variant requires learning only two parameters per dimension, offering a more efficient alternative to more expressive but costly flow parameterizations. For simplicity, we also assume that the initial state prior $p({\mathbf{x}}_{0})$ is Gaussian and known.
@@
The primary objective in ETGPSSM, as in GPSSM, is the joint estimation of model hyperparameters $\bm{\vartheta}\!=\!\{{{\bm{\theta}}}_{gp},\bm{\psi},\mathbf{Q},\mathbf{R}\}$ and inference of both the ETGP and latent states. Given a sequence of observations ${\mathbf{y}}_{1:T}=\{{\mathbf{y}}_{t}\}_{t=1}^{T}$, with $T\in\mathbb{Z}_{+}$, the task is to estimate $\bm{\vartheta}$ and compute the posterior distributions $p(\bm{f}|{\mathbf{y}}_{1:T})$ for the ETGP and $p({\mathbf{x}}_{0:T}|{\mathbf{y}}_{1:T})$ for the latent states, where ${\mathbf{x}}_{0:T}=\{{\mathbf{x}}_{t}\}_{t=0}^{T}$.
However, the marginal likelihood $p(\mathbf{y}_{1:T}|\bm{\vartheta})$ is generally intractable, making both model evidence maximization and posterior inference computationally challenging. Furthermore, unlike standard GPs, which admit closed-form expressions for their prior and posterior (see Eqs. (2) and (3)), the ETGP defines an implicit stochastic process. As a result, both the marginal prior [^4] Obtaining $p(\bm{f})$ by marginalizing over the GP and $\mathbf{w}$ is generally intractable. and posterior distributions are analytically intractable [43], presenting significant challenges for inference. In the following subsections, we introduce a variational inference framework that enables joint learning of model hyperparameters and approximate Bayesian inference over the ETGP and latent states.
@@ III-B Approximate Bayesian Inference
To address the intractability of the marginal distribution $p({\mathbf{y}}_{1:T}|\bm{\vartheta})$, we employ variational inference. This approach maximizes a tractable evidence lower bound (ELBO) on the log-marginal likelihood, which is equivalent to minimizing the Kullback–Leibler (KL) divergence between the variational approximation and the true posterior [49]. Constructing the ELBO requires specifying both the model joint distribution and the variational distribution.
Following the existing GPSSM framework and building upon Eq. (16), we factorize the joint distribution as:
$\!\!\!p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0:T},\bm{f})$ $\!=\!p({\mathbf{y}}_{1:T},{\mathbf{x}}_{1:T}|\bm{f},{\mathbf{x}}_{0})p(\bm{f})p({\mathbf{x}}_{0})$ (17)
$\!=\!\prod_{t=1}^{T}p({\mathbf{y}}_{t}|{\mathbf{x}}_{t})p({\mathbf{x}}_{t}|{\mathbf{x}}_{t-1},\bm{f})p(\bm{f})p({\mathbf{x}}_{0}),$
where $p({\mathbf{x}}_{0})={\cal N}({\mathbf{x}}_{0}|\bm{0},\bm{I})$ and $p(\bm{f})$ denotes the ETGP prior (under slight abuse of notation). A common variational family in the GPSSM literature takes the form:
$q({\mathbf{x}}_{0:T},\bm{f})=q({\mathbf{x}}_{1:T}|\bm{f},{\mathbf{x}}_{0})q(\bm{f})q({\mathbf{x}}_{0}),$ (18)
which yields the ELBO, denoted by $\mathcal{L}_{2}$:
$\mathcal{L}_{2}=\mathbb{E}_{q({\mathbf{x}}_{0:T},\bm{f})}\left[\log\frac{p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0:T},\bm{f})}{q({\mathbf{x}}_{0:T},\bm{f})}\right].$ (19)
While general, this formulation requires explicit parameterization of the variational distribution $q({\mathbf{x}}_{1:T}|\bm{f},{\mathbf{x}}_{0})$, which can introduce a large number of parameters and increase optimization complexity [37, 31].
To improve learning efficiency, in this work, we instead adopt an alternative approach based on marginalizing out ${\mathbf{x}}_{1:T}$ and infer latent states using an EnKF, avoiding the need to explicitly model this variational term. Specifically, marginalizing over ${\mathbf{x}}_{1:T}$ leads to the following joint distribution:
$\!\!\!p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0},\bm{f})$ $\!=\!p({\mathbf{x}}_{0})\,p(\bm{f})\int p({\mathbf{y}}_{1:T},{\mathbf{x}}_{1:T}|\bm{f},{\mathbf{x}}_{0})\,\mathrm{d}{\mathbf{x}}_{1:T}$
$\!=\!p({\mathbf{x}}_{0})\,p(\bm{f})\,p({\mathbf{y}}_{1:T}|{\mathbf{x}}_{0},\bm{f}),$ (20)
where $p({\mathbf{y}}_{1:T}|{\mathbf{x}}_{0},\bm{f})$ will be approximated sequentially using an EnKF, as detailed in Section III-C. Since only ${\mathbf{x}}_{0}$ and $\bm{f}$ remain latent, we adopt the factorized variational distribution:
$q({\mathbf{x}}_{0},\bm{f})=q({\mathbf{x}}_{0})\,q(\bm{f}).$ (21)
This leads to the marginalized ELBO:
$\mathcal{L}=\mathbb{E}_{q({\mathbf{x}}_{0},\bm{f})}\left[\log\frac{p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0},\bm{f})}{q({\mathbf{x}}_{0})\,q(\bm{f})}\right].$ (22)
Compared to $\mathcal{L}_{2}$, the alternative bound $\mathcal{L}$ avoids direct parameterization of $q({\mathbf{x}}_{1:T}\mid\bm{f},{\mathbf{x}}_{0})$, helping reduce optimization overhead and yielding a tighter variational lower bound (see Proposition 1).
Proposition 1.
The bound $\mathcal{L}$ in Eq. (22) satisfies $\mathcal{L}\geq\mathcal{L}_{2}$. Equality holds if and only if the variational approximation $q({\mathbf{x}}_{1:T}|\bm{f},{\mathbf{x}}_{0})$ matches the true posterior $p({\mathbf{x}}_{1:T}|\bm{f},{\mathbf{x}}_{0},{\mathbf{y}}_{1:T})$.
Proof.
Proof can be found in Appendix A
@@
Theoretically, this tighter lower bound can improve inference accuracy and convergence [49]. We now proceed to derive the ELBO $\mathcal{L}$. As noted earlier, the lack of closed-form expressions for the ETGP’s prior and posterior prevents direct computation of $\mathcal{L}$ in function space. To overcome this, we leverage the generative construction of the ETGP—illustrated in Fig. 3—and infer the posterior distributions over $\tilde{f}$ and $\mathbf{w}$, which collectively characterize the ETGP posterior. Specifically, by explicitly representing $\tilde{f}$ and $\mathbf{w}$, the marginalized joint distribution becomes:
$p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0},\mathbf{w},\tilde{f})=p({\mathbf{y}}_{1:T}|{\mathbf{x}}_{0},\mathbf{w},\tilde{f})p({\mathbf{x}}_{0})p_{\bm{\psi}}(\mathbf{w})p(\tilde{f}),$ (23)
where $p(\tilde{f})$ represents the GP prior and $p_{\bm{\psi}}(\mathbf{w})={\cal N}(\mathbf{w}|\bm{0},\operatorname{diag}(\bm{\psi}))$ is a Gaussian prior over the neural network weights, assuming a Bayesian neural network with diagonal covariance. Similarly, instead of using Eq. (21), we employ the following variational distribution:
$q({\mathbf{x}}_{0},\mathbf{w},\tilde{f})=q(\tilde{f})q(\mathbf{w})q({\mathbf{x}}_{0}),$ (24)
which assumes independence between $\tilde{f}$, $\mathbf{w}$ and ${\mathbf{x}}_{0}$. Here the variational distributions of ${\mathbf{x}}_{0}$ and $\mathbf{w}$ are defined as follows:
[…]
With Eqs. (23) and (24), we can derive our ELBO, $\mathcal{L}$, which is summarized in the following proposition.
Proposition 2.
Under all the assumed variational distributions, together with the model joint distribution in Eq. (23), the ELBO can be reformulated as follows:
$\mathcal{L}=$ $\mathbb{E}_{q(\mathbf{w},\tilde{f})}\left[\sum_{t=1}^{T}\log p({\mathbf{y}}_{t}|{\mathbf{y}}_{1:t-1},\mathbf{w},\tilde{f})\right]-\operatorname{KL}(q({\mathbf{u}})\|p({\mathbf{u}}))$
$-\operatorname{KL}(q(\mathbf{w})\|p(\mathbf{w}))-\operatorname{KL}(q({\mathbf{x}}_{0})\|p({\mathbf{x}}_{0})),$ (26)
where the first terms (log-likelihood) can be analytically evaluated using the EnKF (detailed in Section III-C). The three KL divergence terms can also be computed in closed form, due to the Gaussian nature of the prior and variational distributions [49].
Proof.
The proof can be found in Appendix B. ∎
Remark 2.
The objective in Eq. (26) can also be derived from $\mathcal{L}_{2}$ under two specific assumptions (see Supplement A), akin to [37]. However, our derivation in this work is generic for arbitrary random processes and is assumption-free, offering greater simplicity, conceptual clarity, and broader applicability. For further details on the relationship between $\mathcal{L}$ and $\mathcal{L}_{2}$, we refer to Supplement  A.
III-C EnKF-aided ELBO Evaluation
This section describes how to evaluate $p({\mathbf{y}}_{t}|{\mathbf{y}}_{1:t-1},\mathbf{w},\tilde{f})$ in Eq. (26) using the EnKF. First, note that
$p({\mathbf{y}}_{t}|{\mathbf{y}}_{1:t-1},\mathbf{w},\tilde{f})$ $=\int\underbracket{p({\mathbf{y}}_{t}|{\mathbf{x}}_{t})}_{\text{likelihood}}\underbracket{p({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t-1})}_{\text{prediction}}\mathrm{d}{\mathbf{x}}_{t},$ (27)
where the prediction distribution $p({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t-1})$ reads:
The transition term can be expressed as